RADIATIVE LIFETIMES OF NiH AND CoH

\underline{JENNIFER GOTTFRIED} and JEFFREY A. GRAY, Department of Chemistry, Ohio Northern University, Ada, OH 45810.

Radiative lifetimes have been measured for several excited electronic states of NiH and CoH radicals, which are produced by laser ablation in a pulsed free-jet expansion of argon containing pentane vapor. Known^{a}^{b} transitions in the 15,500 to 18,200 cm^{−1} region were excited using a pulsed dye laser, and lifetimes varying from approx. 0.5 to 5 μs were determined by temporally resolving the fluorescence intensity. The lifetimes of some states depend on rotational level and/or parity. The electronic structure of transition-metal radicals is generally difficult to model or theoretically predict, especially for optically excited states. While observations such as magnetic moments, Λ-doubling, and spin-orbit intervals can aid in understanding perturbations, radiative lifetimes may be a better diagnostic for assigning the spin multiplicity of the excited states when the low-lying states have well-defined spin.

aS. A. Kadavathu, R. Scullman, R. W. Field, J. A. Gray, and M. Li, J. Mol. Spectrosc. \underline{147}, 448-470 (1991).

bT. D. Varberg, E. J. Hill, and R. W. Field, J. Mol. Spectrosc. \underline{138}, 630-637 (1989).